Understanding a particular spectrum is important i
Answer #1

\lambda = 590 nm ,n =1, \theta =34.7 degrees

dsin\theta = n\lambda

dsin(34.7) = 1*590*10^-9

d = 1036.4 nm

(b) dsin\theta = n\lambda

for \lambda = 656.3 nm

1036.4*10^-9 sin\theta =1*656.3*10^-9

\theta = 39.3 degrees

for \lambda = 486.1 m

1036.4*10^-9 sin\theta =1*486.1*10^-9

\theta = 27.97 degrees